# Fine-Tuning Diffusion Generative Models via Rich Preference Optimization

### v4

@@ Abstract
We introduce Rich Preference Optimization (RPO), a novel pipeline that leverages rich feedback signals to improve the curation of preference pairs for fine-tuning text-to-image diffusion models. Traditional methods, like Diffusion-DPO, often rely solely on reward model labeling, which can be opaque, offer limited insights into the rationale behind preferences, and are prone to issues such as reward hacking or overfitting. In contrast, our approach begins with generating detailed critiques of synthesized images, from which we extract reliable and actionable image editing instructions. By implementing these instructions, we create refined images, resulting in synthetic, informative preference pairs that serve as enhanced tuning datasets. We demonstrate the effectiveness of our pipeline and the resulting datasets in fine-tuning state-of-the-art diffusion models. Our code is available at https://github.com/Diffusion-RLHF/RPO.
1Department of IEOR, Columbia University 2Amazon
3Department of ORFE, Princeton University 4Capital One
11footnotetext: Equal contribution.22footnotetext: Work done outside Amazon.33footnotetext: Work done outside Capital One.
@@ 1 Introduction
Learning from feedback and critiques is essential for enhancing the performance of a generative model by guiding the model to rectify unsatisfactory outputs, e.g. in RLHF (Ouyang et al., 2022). However, improvements can arise not only from distinguishing right from wrong, but also from receiving thoughtful and informative feedback that offers clear direction for enhancement, which we refer as rich feedback. For instance, in the natural language tasks, rich feedback has proved to be useful in LLM self-refinement (Saunders et al., 2022), reward model enhancement (Ye et al., 2024; Ankner et al., 2024), code debugging (Chen et al., 2023; McAleese et al., 2024), games (Paglieri et al., 2024; Hudi et al., 2025), and agents (Shinn et al., 2023; Wang et al., 2023; Renze & Guven, 2024). Rich feedback is comparatively less explored in vision tasks, such as human feedback generation model (Liang et al., 2024) and visual commonsense reasoning (Cheng et al., 2024; Chen et al., 2024; Li et al., 2024c).
@@
To leverage rich feedback in preference learning, we draw inspiration from the way students learn from their teachers. We introduce Rich Preference Optimization (RPO), a novel approach designed to enhance preference learning for images by leveraging vision-language models (VLMs). These multi-modal models provide detailed critiques for downstream tasks need, offering rich feedback that mirrors the comprehensive guidance students receive in modern educational systems. Rather than merely receiving a final score on assignments or exams, students are given specific feedback that identifies logical errors, misunderstandings, or calculation mistakes. This feedback enables students to iteratively refine their initial responses, fostering deeper learning through a process of continuous improvement. Similarly, in RPO, we extract actionable editing instructions from VLMs and employ instruction-driven image-editing models for refinement. This scalable method generates informative preference pairs that are crucial for effective preference learning. The process of receiving detailed feedback and making refinements is akin to learning from true preference pairs, reflecting the natural and effective way in which humans learn.
Figure 1: (Top) Our RPO pipeline for curating informative preference pairs from images generated from the base diffusion models: (1) Rich Feedback/Critic generation by a Vision Language Model (for which we choose LLaVA-Critic-7B), (2) Actionable editing instruction generation based on the critiques by another VLM (for which we chose Qwen2.5-VL-8B-Instruct), (3) Instruction-following image editing from the generated editing instructions (for which we choose ControlNet), and (4) Diffusion DPO training using reward model filtered synthetic preference pairs. (Bottom) Sample images generated from RPO fine-tuned Stable Diffusion XL, by further aligning the model on our generated synthetic preferences.
The contribution of our paper is summarized as follows:
@@
    We also empirically demonstrate that our curated preferences are scalable and useful offline synthetic preference dataset, by further training Diffusion DPO (trained on the original offline dataset) checkpoints on rich preferences we generate on the preferred images. We observe drastic performance gain, which validates the effectiveness of our pipeline. We include qualitative samples of our obtained checkpoints also in Figure 1.
@@
Concurrent work to ours, (Wang et al., 2025) also investigates on utilizing dedicated feedback and then passing to the editing model in the contents of an inference time scaling method. As a comparison, we focus on alignment and our pipeline has an extra procedure of generating concrete editing instructions before passing to instruction-following editing models, which directly connects two well developed literature of VLMs and Image-editing models. Our pipeline is also easier for plug-in thanks to the recent rapid progress in instruction-following image-editing capability like in GPT-4o (Hurst et al., 2024). Our work is also reminiscent to Aligner (Ji et al., 2024), which directly tries to find a more preferred answer.
@@ 2 Preliminaries
Diffusion Models. Diffusion models are a class of generative models $p_{\theta}(x_{0})$, whose goal is to turn noisy/non-informative initial distribution $p_{\tiny\mbox{noise}}(x_{T})$ to a desired target distribution $p_{\tiny\mbox{tar}}(x_{0})$ through a well-designed denoising process (Ho et al., 2020; Song et al., 2021a; b). Here we adopt the discrete-time formulation of diffusion models as in Ho et al. (2020).
Given noise scheduling functions $\alpha_{t}$ and $\sigma_{t}$ (as defined in Rombach et al. (2022)), the forward process is specified by $q(x_{t}\,|\,x_{s})=\mathcal{N}\left(\frac{\alpha_{t}}{\alpha_{s}}x_{s},\,\left(\sigma_{t}^{2}-\frac{\alpha_{t}^{2}}{\alpha_{s}^{2}}\sigma_{s}^{2}\right)I\right)$ for $s<t$. Its time-reversed process is a Markov chain parameterized by $p_{\theta}(x_{0:T})=\prod_{t=1}^{T}p_{\theta}(x_{t-1}\,|\,x_{t})\,p_{\tiny\mbox{noise}}(x_{T})$, where
$p_{\theta}(x_{t-1}\,|\,x_{t})=\mathcal{N}\left(\mu_{\theta}(t,x_{t}),\,\Sigma_{t}I\right),$ (1)
with $\mu_{\theta}(t,x):=\frac{\alpha_{t-1}}{\alpha_{t}}x_{t}-(\frac{\alpha_{t-1}}{\alpha_{t}}-\frac{\alpha_{t}\sigma_{t-1}^{2}}{\alpha_{t-1}\sigma_{t}^{2}})\sigma_{t}\epsilon_{\theta}(t,x_{t})$ and $\Sigma_{t}=\sigma_{t-1}^{2}-\frac{\alpha_{t}^{2}\sigma_{t-1}^{4}}{\alpha_{t-1}^{2}\sigma_{t}^{2}}$. The model equation 1 is trained by minimizing the evidence lower bound (ELBO):
$\min_{\theta}\mathcal{L}(\theta):=\mathbb{E}_{t,\epsilon,x_{0},x_{t}}\left[\omega\left(\lambda_{t}\right)||\epsilon-\epsilon_{\theta}(t,x_{t})||_{2}^{2}\right],$ (2)
where $t\sim\mathcal{U}(0,T)$, $\epsilon\sim\mathcal{N}(0,I)$, $x_{0}\sim p_{\tiny\mbox{tar}}(x_{0})$, $x_{t}\sim q(x_{t}\,|\,x_{0})=\mathcal{N}(\alpha_{t}x_{0},\,\sigma_{t}^{2}I)$, $\lambda_{t}:=\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}}$ is the signal-to-noise ratio, and $\omega:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is some weight function. The training process equation 2 is also known as denoising score matching (see Vincent (2011) or (Tang & Zhao, 2024, Section 4.3)). It is expected that for $\theta_{*}$ solving the optimization problem equation 2, the model’s output distribution $p_{\theta_{*}}(x_{0})\approx p_{\tiny\mbox{tar}}(x_{0})$, see Chen et al. (2022); Li et al. (2024a) for the theory.
@@
As previously mentioned, we edit images based on ControlNet (Zhang et al., 2023b) to ensure a better coherence to the original image. ControlNet has been widely used for controlling image diffusion models by conditioning the model with an additional input image. Further applications include implementations for the state-of-the-art proprietary models like Stable Diffusion 3, Stable Diffusion XL (Esser et al., 2024) and FLUX (Labs, 2024), multi-image support extensions like ControlNet++ (Li et al., 2024b). We will stick to the original ControlNet in this paper (because the offline dataset is generated by the similar model scaled by Stable Diffusion v1.5), and leave the usage of more advanced ControlNet variants in future work.
@@
$L_{\mathrm{DPO}}(\theta)=-\mathbb{E}_{c,x_{0}^{w},x_{0}^{l}}\log\sigma\Biggl{(}\beta\mathbb{E}_{\begin{subarray}{c}x_{1:T}^{w}\sim p_{\theta}(x_{1:T}^{w}|x_{0}^{w},c)\\ x_{1:T}^{l}\sim p_{\theta}(x_{1:T}^{l}|x_{0}^{l},c)\end{subarray}}\bigg{[}\log\frac{p_{\theta}(x_{0:T}^{w}|c)}{p_{\mathrm{ref}}(x_{0:T}^{w}|c)}-\log\frac{p_{\theta}(x_{0:T}^{l}|c)}{p_{\mathrm{ref}}(x_{0:T}^{l}|c)}\bigg{]}\Biggr{)}.$ (3)
By Jensen’s inequality and approximate $p_{\theta}(x_{1:T}|x_{0},c)$ by the forward process $q(x_{1:T}|x_{0})$, Wallace et al. (2024) yields the following loss function for efficient training:
$L_{\mathrm{Diffusion-DPO}}(\theta)=-\mathbb{E}_{x_{0}^{w},x_{0}^{l},t,x_{t}^{w}\sim q(x_{t}^{w}|x_{0}^{w}),x_{t}^{l}\sim q(x_{t}^{l}|x_{0}^{l})}\log\sigma\biggl{(}\beta T\omega(\lambda_{t})$ (4)
$\Big{(}-\|\epsilon^{w}-\epsilon_{\theta}(x_{t}^{w},t)\|_{2}^{2}+\|\epsilon^{w}-\epsilon_{\mathrm{ref}}(x_{t}^{w},t)\|_{2}^{2}+(\|\epsilon^{l}-\epsilon_{\theta}(x_{t}^{l},t)\|_{2}^{2}-\|\epsilon^{l}-\epsilon_{\mathrm{ref}}(x_{t}^{l},t)\|_{2}^{2})\Big{)}\biggr{)},$
where $x_{t}^{w,l}\sim q(x_{t}^{w,l}\,|\,x_{0}^{w,l})=\mathcal{N}(\alpha_{t}x_{0}^{w,l}\sigma_{t}^{2}I)$, and $\epsilon^{w,l}\in\mathcal{N}(0,I)$.
@@ 3.1 Instruction-Following Image Editing
To tackle the above issue, we propose a ControlNet (Zhang et al., 2023b) based image-editing approach. Concretely, we exploit the image2image (pipeline) of the ControlNet by setting the conditional image to be the same as the original one, which guarantees that the changes adhere to the original image. In addition, we concatenate the prompt that generates the image with our generated editing instructions, which we find will greatly enhance the quality of the edited image generation as illustrated in Figure 2.
@@ 3.2 Generating Rich Feedback Signals
To compare and ablate about what type of feedback will be more useful for yielding better editing instructions for improving the images, we prompt GPT-4o to generate them based on the given feedback, and then compute the rewards of new images (evaluated by different reward models) after using ControlNet to edit the image with the instructions as in Figure 4(a). Concretely, we let GPT-4o generate editing instructions based on various types of feedback, including those from RichFB, LLaVA-Critic, and even ChatGPT-4o itself. From the RichFB dataset, we use the image, prompt, misalignment information within the prompt, and the misalignment heatmap of the image as inputs. From LLaVA-Critic, we incorporate its textual feedback. For ChatGPT-4o, we first prompt it to also generate textual feedback based on the image-prompt pair, which is then used to derive editing instructions. For studing the necessity of rich feedback, we also let GPT 4o generate editing instructions directly from the input image and prompt, bypassing the intermediate step of critique.
@@ 4.1 Settings
Diffusion-DPO Training. For Diffusion-DPO training, we follow the same setting and use the same hyperparameters in Wallace et al. (2024) when fine-tuning SD1.5-based models. More specifically, we use AdamW (Loshchilov & Hutter, 2019) as the optimizer; the effective batch size is set to 2048; we train at fixed square resolutions, and use a learning rate of $\frac{2000}{\beta}2.048\cdot 10^{-8}$ with $25\%$ linear warmup; for the divergence penalty parameter, we keep $\beta=5000$. When fine-tuning SDXL-based models, we use the LoRA implementation provided by the Github project (Suzukimain & Liu, 2024) to improve training efficiency.
@@
   DPO-SD1.5-100k (ImageReward-Aligned) 20.45 0.3913 6.097 0.2645
   DPO-SD1.5-100k (ImageReward-Aligned) + RPO100k 20.54 0.5252 6.105 0.2660
@@ 4.2 Improving Diffusion-DPO via RPO
We present the performance of all models evaluated by different reward models in Table 1. Above all, models obtained by fine-tuning on our 100k synthetic preferences data achieve comparable levels of the baseline models by all metrics and outperform the corresponding baseline models in most cases. This confirms our hypothesis that the pipeline we designed for synthesizing preference pairs yields improved results when combined with preference-based training algorithms such as Diffusion-DPO.
@@
Comparing the fine-tuned model DPO-SD1.5-100k+RPO100k with the baseline model DPO-SD1.5-200k, we note that our pipeline leads to a higher data efficiency for Diffusion-DPO training, in the sense that the former model utilizes a mixture of 100k human-labeled data and 100k synthetic data, while the latter uses 200k human-labeled data, which is naturally considered to be of higher quality. We have thus verified the effectiveness and potential of the synthetic preferences data generated by rich feedback to augment the given (high-quality) dataset, which enhances the performance of fine-tuning. This is of crucial importance when human annotationed data is limited (which is usually always the case) .
@@
We also present a qualitative comparison of the three SDXL-based models in Figure 6. In addition to generating images with better visual qualities, we remark that the RPO fine-tuned model is better at connecting different elements in the prompts in a deep and profound way (see the first and the third columns). Understanding how rich feedback and guided revision help fine-tuning diffusion models will be left for future work.
4.3 Data Scaling Analysis and Generalization to Diverse Prompt Distributions
To further evaluate the robustness and scalability of our method, we conduct additional experiments on data scaling. Specifically, we subsample the vanilla preference dataset used in Diffusion-DPO training at sizes of 5K, 10K, 20K, 50K, and 100K, and compare performance between Diffusion-DPO and our Diffusion-DPO+RPO pipeline. The metrics of full Diffusion-DPO (denoted as star in Figure 7) are reported based on the evaluation of checkpoints released by Wallace et al. (2024), which is trained on nearly 1M offline preference pairs.
Scalability. As shown in all plots in Figure 7, DPO + RPO consistently improves upon DPO across all preference set sizes, with gains becoming increasingly pronounced as the number of training data increases. This trend holds across multiple evaluation metrics—including PickScore, ImageReward, LAION-Aesthetic, and style-specific scores (Photo, Paintings, Anime, and Concept Art), and across prompt sets such as Pick-a-Pic, PartiPrompts, and HPSv2.
Data efficiency. Notably, our DPO-SD1.5 model fine-tuned with only 100K vanilla preferences and enriched with 100K rich-feedback pairs (DPO-SD1.5-100k + RPO100k) outperforms the Diffusion-DPO model trained on nearly 1M Pick-a-Pic preferences on all major metrics except LAION-Aesthetic. In contrast, we observe the full DPO model exhibits inconsistent performance across reward metrics (e.g., similar ImageReward scores on Pick-a-pic and PartiPrompts and lower LAION-Aesthetic score on PartiPrompts, and more on HPSv2), suggesting overfitting to narrow preference signals. These findings highlight the strong data efficiency and generalization capabilities of RPO: it achieves superior or comparable performance using an order of magnitude less training data, without requiring additional inference or human labeling to generate new data.
Generalization. We further examine the generalization behavior of RPO using two out-of-distribution benchmarks: PartiPrompts and HPSv2. The PartiPrompts set evaluates transfer to novel prompt distributions, while the HPSv2 benchmark assesses robustness across diverse visual styles, including Photo, Paintings, Anime, and Concept Art. As shown in Figure 7, RPO-enhanced models consistently outperform the Diffusion-DPO baseline across all style types, prompt sets, and dataset sizes. The improvements are especially notable in the Paintings and Concept Art categories, where DPO+RPO exhibits steeper and more consistent gains as the number of preference pairs increases. The average HPSv2 score reflects this trend, with DPO+RPO maintaining a significant margin over DPO at every scale. These results highlight RPO’s robustness in handling both stylistic and prompt-based distribution shifts, confirming that the benefits of rich feedback generalize effectively across domains.

### v5

# RPO: Fine-Tuning Visual Generative Models via Rich Vision-Language Preferences

@@ Abstract
Traditional preference tuning methods for LLMs/Visual Generative Models often rely solely on reward model labeling, which can be opaque, offer limited insights into the rationale behind preferences, and are prone to issues such as reward hacking or overfitting. We introduce Rich Preference Optimization (RPO), a novel pipeline that leverages rich feedback signals from Vision Language Models (VLMs) to improve the curation of preference pairs for fine-tuning visual generative models like text-to-image diffusion models. Our approach begins with prompting VLMs to generate detailed critiques of synthesized images, from which we further prompt VLMs to extract reliable and actionable image editing instructions. By implementing these instructions, we create refined images, resulting in synthetic, informative preference pairs that serve as enhanced tuning datasets. We demonstrate the effectiveness of our pipeline and the resulting datasets in fine-tuning state-of-the-art diffusion models.
RPO: Fine-Tuning Visual Generative Models via Rich Vision-Language Preferences
[…]
Tingting Ou1 1Columbia University          Ziyu Huang1 1Columbia University          David D. Yao1 1Columbia University          Wenpin Tang1 1Columbia University
Figure 1: Our RPO pipeline for curating informative preference pairs from images generated from the base diffusion models: (1) Rich Feedback/Critic generation by a VLM (for which we choose LLaVA-Critic-7B), (2) Actionable editing instruction generation based on the critiques by another VLM (for which we chose Qwen2.5-VL-8B-Instruct), (3) Instruction-following image editing from the generated editing instructions (for which we choose ControlNet), and (4) Diffusion DPO training using reward model filtered preference pairs.
@@ 1 Introduction
Learning from feedback and critiques is essential for enhancing the performance of a generative model by guiding the model to rectify unsatisfactory outputs, e.g. in RLHF (Ouyang et al., 2022). However, improvements can arise not only from distinguishing right from wrong, but also from receiving thoughtful and informative feedback that offers clear direction for enhancement, which we refer as rich feedback. For instance, in the natural language tasks, rich feedback has proved to be useful in LLM self-refinement (Saunders et al., 2022), reward model enhancement (Ye et al., 2024; Ankner et al., 2024), code debugging (Chen et al., 2023; McAleese et al., 2024), games (Paglieri et al., 2024; Hudi et al., 2025), and agents (Shinn et al., 2023; Wang et al., 2023; Renze and Guven, 2024). Rich feedback is comparatively less explored in vision tasks, such as human feedback generation model (Liang et al., 2024) and visual commonsense reasoning (Cheng et al., 2024; Chen et al., 2024; Li et al., 2024b).
@@
To leverage rich feedback in preference learning, we draw inspiration from the way students learn from their teachers. We introduce Rich Preference Optimization (RPO), a novel approach designed to enhance preference learning of images by letting the model learn from rich preferences that we curate with the aid of vision-language models (VLMs). These multi-modal models provide detailed critiques for downstream tasks need, offering rich feedback that mirrors the comprehensive guidance students receive in modern educational systems. Rather than merely receiving a final score on assignments or exams, students are given specific feedback that identifies logical errors, misunderstandings, or calculation mistakes. This feedback enables students to iteratively refine their initial responses, fostering deeper learning through a process of continuous improvement. Similarly, in RPO, we extract actionable editing instructions from VLMs and employ instruction-driven image-editing models for refinement. This scalable method generates informative preference pairs that are crucial for effective preference learning. The process of receiving detailed feedback and making refinements is akin to learning from true preference pairs, reflecting the natural and effective way in which humans learn. The contribution of our paper is summarized as follows:
@@
    We also empirically demonstrate that our curated preferences are scalable and useful offline synthetic preference dataset, by further training Diffusion DPO (trained on the original offline dataset) checkpoints on rich preferences we generate on the preferred images. We observe drastic performance gain, which validates the effectiveness of our pipeline.
@@
Concurrent work to ours, Wang et al. (2025) also investigate on utilizing dedicated feedback and then passing to the editing model in the contents of an inference time scaling method. As a comparison, we focus on alignment and our pipeline has an extra procedure of generating concrete editing instructions before passing to instruction-following editing models, which directly connects two well developed literature of VLMs and Image-editing models. Our pipeline is also easier for plug-in thanks to the recent rapid progress in instruction-following image-editing capability like in GPT-4o (Hurst et al., 2024). Our work is also reminiscent to Aligner (Ji et al., 2024) for pure language setup.
@@ 2 Preliminaries
We discussion some preliminaries about visual generative models like diffusion models, VLMs and relevant image-editing models.
Diffusion Models. Diffusion models are a class of generative models $p_{\theta}(x_{0})$, whose goal is to turn noisy/non-informative initial distribution $p_{\mbox{noise}}(x_{T})$ to a desired target distribution $p_{\mbox{tar}}(x_{0})$ through a well-designed denoising process (Ho et al., 2020; Song et al., 2021a, b). Given noise scheduling functions $\alpha_{t}$ and $\sigma_{t}$ (as defined in Rombach et al. (2022)), the forward process is specified by $q(x_{t}\,|\,x_{s})=\mathcal{N}\left(\frac{\alpha_{t}}{\alpha_{s}}x_{s},\,\left(\sigma_{t}^{2}-\frac{\alpha_{t}^{2}}{\alpha_{s}^{2}}\sigma_{s}^{2}\right)I\right)$ for $s<t$. A diffusion model is trained by minimizing the evidence lower bound (ELBO):
[…]
where $t\sim\mathcal{U}(0,T)$, $\epsilon\sim\mathcal{N}(0,I)$, $x_{0}\sim p_{\mbox{tar}}(x_{0})$, $x_{t}\sim q(x_{t}\,|\,x_{0})=\mathcal{N}(\alpha_{t}x_{0},\,\sigma_{t}^{2}I)$, $\lambda_{t}:=\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}}$ is the signal-to-noise ratio, and $\omega:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is some weight function.
@@
As previously mentioned, we edit images based on ControlNet (Zhang et al., 2023b) to ensure a better coherence to the original image. ControlNet has been widely used for controlling image diffusion models by conditioning the model with an additional input image. Further applications include implementations for the state-of-the-art proprietary models like Stable Diffusion 3, Stable Diffusion XL (Esser et al., 2024) and FLUX (Labs, 2024), multi-image support extensions like ControlNet++ (Li et al., 2024a). We will stick to the original ControlNet in this paper (because the offline dataset is generated by the similar model scaled by Stable Diffusion v1.5), and leave the usage of more advanced ControlNet variants in future work.
@@
$L_{\mathrm{DPO}}(\theta)$ $=-\mathbb{E}_{c,x_{0}^{w},x_{0}^{l}}\log\sigma(\beta\mathbb{E}_{\begin{subarray}{c}x_{1:T}^{w}\sim p_{\theta}(\cdot|x_{0}^{w},c)\\ x_{1:T}^{l}\sim p_{\theta}(\cdot|x_{0}^{l},c)\end{subarray}}$ (2)
$\qquad\bigg[\log\frac{p_{\theta}(x_{0:T}^{w}|c)}{p_{\mathrm{ref}}(x_{0:T}^{w}|c)}-\log\frac{p_{\theta}(x_{0:T}^{l}|c)}{p_{\mathrm{ref}}(x_{0:T}^{l}|c)}\bigg])$
@@ 3.1 Instruction-Following Image Editing
To tackle the above issue, we propose a ControlNet (Zhang et al., 2023b) based image-editing approach. Concretely, we exploit the image2image (pipeline) of the ControlNet by setting the conditional image to be the same as the original one, which guarantees that the changes adhere to the original image. In addition, we concatenate the prompt that generates the image with our generated editing instructions, which we find will greatly enhance the quality of the edited image generation as illustrated in Appendix B.
@@ 3.2 Generating Rich Feedback Signals
To compare and ablate about what type of feedback will be more useful for yielding better editing instructions for improving the images, we prompt GPT-4o to generate them based on the given feedback, and then compute the rewards of new images (evaluated by different reward models) after using ControlNet to edit the image with the instructions as in Figure 3(a). Concretely, we let GPT-4o generate editing instructions based on various types of feedback, including those from RichFB, LLaVA-Critic, and even GPT-4o itself. From the RichFB dataset, we use the image, prompt, misalignment information within the prompt, and the misalignment heatmap of the image as inputs. From LLaVA-Critic, we incorporate its textual feedback. For GPT-4o, we first prompt it to also generate textual feedback based on the image-prompt pair, which is then used to derive editing instructions. For studing the necessity of rich feedback, we also let GPT 4o generate editing instructions directly from the input image and prompt, bypassing the intermediate step of critique.
@@ 4.1 Settings
Diffusion-DPO Training. For Diffusion-DPO training, we follow the same setting and use the same hyperparameters in Wallace et al. (2024) when fine-tuning SD1.5-based models. More specifically, we use AdamW (Loshchilov and Hutter, 2019) as the optimizer; the effective batch size is set to 2048; we train at fixed square resolutions, and use a learning rate of $\frac{2000}{\beta}2.048\cdot 10^{-8}$ with $25\%$ linear warmup; for the divergence penalty parameter, we keep $\beta=5000$. When fine-tuning SDXL-based models, we use the LoRA implementation provided by the Github project (Suzukimain and Liu, 2024) to improve training efficiency.
@@
 DPO-SD1.5-100k (ImageReward-Aligned) 20.66 0.600 6.083 0.268
 DPO-SD1.5-200k (ImageReward-Aligned) 20.63 0.606 6.109 0.268
 DPO-SD1.5-100k (ImageReward-Aligned) + RPO100k 20.72 0.686 6.213 0.269
@@ 4.2 Improving Diffusion-DPO via RPO
We present the performance of all models evaluated by different reward models in Table 1. Above all, models obtained by fine-tuning on our 100k synthetic preferences data outperforms the baseline models by all metrics. This confirms our hypothesis that the pipeline we designed for synthesizing preference pairs yields improved results when combined with preference-based training algorithms such as Diffusion-DPO.
@@
Concretely, when utilizing the same amount of vanilla preference data (synthetic data for RPO is curated from this same vanilla preference data), our RPO pipeline stably improves the DPO baseline over all metrics, see DPO-SD1.5-100k and DPO-SD1.5-100k+RPO100k. Even comparing with the model DPO-SD1.5-200k which utilized twice the additional vanilla preference data, our RPO fine-tuned model still performs better across all metrics. This demonstrates that our pipeline is significantly more data efficient for Diffusion-DPO style of training than just on vanilla preference data, achieving more than 100% gain over the standard practice. We have thus verified the effectiveness and potential of the synthetic preferences data generated by rich feedback to augment the given (high-quality) dataset, which is of crucial importance when human annotated data is limited (which we believe is usually always the case in practice).
@@
We also present a qualitative comparison of the three SDXL-based models in Appendix Sec C.1 (see Figure 8). Beyond generating images with better visual qualities, we remark that the RPO fine-tuned model is better at connecting different elements in the prompts in a deep and profound way (see the first and the third columns). Understanding how rich feedback and guided revision help fine-tuning diffusion models will be left for future work.
4.3 Data Scaling and Generalization Analysis
We evaluate the robustness of our method by comparing performance across training scales ranging from 5K to 100K samples. As illustrated in Figure 5, RPO-enhanced models consistently outperform the Diffusion-DPO baseline across all preference set sizes, with the performance margin widening as the training volume increases. Most notably, the model fine-tuned on only 100K RPO-enriched pairs outperforms the standard Diffusion-DPO model trained on nearly 1M pairs (denoted by stars in Figure 5) across all major metrics.
This order-of-magnitude improvement in data efficiency highlights RPO’s ability to achieve superior performance without additional human labeling. In contrast, we observe that the 1M-sample baseline exhibits inconsistent performance across reward metrics, such as lower LAION-Aesthetic scores on PartiPrompts despite the increased data volume, suggesting that it overfits to narrow preference signals rather than learning generalized visual alignment. Furthermore, RPO demonstrates strong generalization capabilities on out-of-distribution benchmarks, specifically PartiPrompts and HPSv2 (see additional results in Appendix E, Figure 13). These gains are consistent across diverse visual styles, confirming that the benefits of learning from rich feedback generalize effectively across novel prompt and stylistic distributions.